OpenStudy (anonymous):

Find x in 4^x-9^x = 0

OpenStudy (zpupster):

1-1=0 what x will give you 1

OpenStudy (ikram002p):

@zpupster :)

OpenStudy (anonymous):

huh? I don't get it. Can you explain thoroughly. Please?

OpenStudy (anonymous):

@zpupster

OpenStudy (zpupster):

anything raised to the 0 is 1

OpenStudy (cwrw238):

x = 0 is the answer

OpenStudy (ikram002p):

to solve this now u need to add 9^x to both side 4^x=9^x... now x is the same in both cases but 4 dnt equal 9 so there is no integer number (x)>0 could make this true unless u know the rule of a^0=1,,, such that a any num.

OpenStudy (cwrw238):

zpupster had the right idea except he made a human error ( x = 1 was incorrect)

OpenStudy (anonymous):

so do you mean? 4^x - 9^x = 0 1-1=0 x=0 checking 4^0-9^0 1-1=0 0=0

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

is my solution correct?

OpenStudy (cwrw238):

yes

OpenStudy (ikram002p):

u r correct

OpenStudy (ikram002p):

@cwrw238 zpupster give 0 as an answer too !!

OpenStudy (cwrw238):

x^0 , 4^0 - anything to power 0 = 1

OpenStudy (anonymous):

hmm. Now i see :) hehe. Thank yoou so much for your help guys. I don't know whick one of you i will award a best response

OpenStudy (anonymous):

i have another problem. Can you help me guys? just last two. Please?

OpenStudy (anonymous):

look. find x if 2^(3-x) = 565

OpenStudy (cwrw238):

oops - sorry - i misread his zpupster's first post

OpenStudy (anonymous):

find x if 2^(3-x) = 565

OpenStudy (anonymous):

please :(

OpenStudy (anonymous):

i really can't get the answer.

OpenStudy (anonymous):

did you get the answer? @ikram002p

OpenStudy (ikram002p):

im trying but r u sure from 565 ??

OpenStudy (zpupster):

ok try take reciprocal 2^x-3 = 1/565 then log base 2 both sides x-3 = - log(565)/log(2) add 3 both sides

OpenStudy (anonymous):

yes. it's 565. :3 my geo prof. Told me to use logarithmic, coz 565 is impossible to be in exponential form of 2. But, we are not yet through with logarithmic law

OpenStudy (anonymous):

@zpupster i don't get it. T_T

OpenStudy (ikram002p):

mmm the qs couldnt be solved without log... so u need to learn log to solve it ..

OpenStudy (ikram002p):